OpenStudy (anonymous):

What mixed number that falls in between of 1 4/5 and 2

sammixboo (sammixboo):

Okie dokie can you convert the fraction \(\sf \large \color{magenta}{1\frac{4}{5}}\) to \(\sf \large \color{magenta}{tenths}\)?

sammixboo (sammixboo):

To do so you have to take \(\sf \large \color{magenta}{\frac{4}{5}}\) and multiply both the \(\sf \large \color{magenta}{4}\) and the \(\sf \large \color{magenta}{5}\) and multiply it by \(\sf \large \color{magenta}{2}\) to get the fraction \(\sf \large \color{magenta}{\frac{8}{10}}\) then you just add the whole number \(\sf \large \color{magenta}{1}\) to that fraction, so it will be \(\sf \large \color{magenta}{1\frac{8}{10}}\)

sammixboo (sammixboo):

What fraction is in the middle of \(\sf \large \color{magenta}{1\frac{8}{10}}\) and \(\sf \large \color{magenta}{2}\)

OpenStudy (anonymous):

Thank you so much!!

sammixboo (sammixboo):

You are welcome! You can tag me in any question you have by typing \(\sf \large \color{magenta}{@sammixboo}\) in the comments.

sammixboo (sammixboo):

Like this... @ace120

sammixboo (sammixboo):

And do you see the 'Best Response' button over there ---------------------->

sammixboo (sammixboo):

Click it to ive someone a medal to show that they gave the best response!